Determine the empirical formula if a compound is made up of 14.0 g Nitrogen and 8.00 g Oxygen. 



1
Expert's answer
2022-04-04T17:05:01-0400

N2O

m(N) = 14 g

Mr(N) = 14

m(O) = 8 g

Mr(O) = 16

N : O = n(N)/n(O) = m(N)/Mr(N) : m(O)/Mr(O)

N : O = 14/14 : 8/16 = 1 : 0,5 => *2 => N : O = 2:1 => N2O